OpenStudy (anonymous):

solve the non-standard equation for x, -10 + x=-1 . help plz how do I set up this problem?

OpenStudy (anonymous):

get x on its own add 10 to both sides

OpenStudy (anonymous):

-10+x=-1 +10 +10 x=9